OpenStudy (shamil98):

If b^p = N, then which of the following is true? log_b N = p log_p N = b log_N b =p N^1/b = p Not that familiar with log stuff, could someone assist me?

OpenStudy (phi):

If you have time, see https://www.khanacademy.org/math/algebra/logarithms-tutorial/logarithm_basics/v/logarithms logs "undo" exponents. if you have \[ a^b = c \\ \text{ take the log base a of both sides}\\ \log_a(a^b) =\ \log_a(c) \\ b = \log_a(c) \]

OpenStudy (shamil98):

Oh. I think I understand so. \[b^p = N\] \[\log_b b^p = \log_b N\] \[p = \log_b N\]

OpenStudy (ranga):

Yes.

OpenStudy (shamil98):

Thank you both.